OpenStudy (anonymous):

find an equation of the line having the given slope and containing the give point m=-6, (8,9)

OpenStudy (anonymous):

Use the formula y=mx+b Plug in what you have \[9=-6\times8+b\] solve for b \[9 = -48 +b \] \[57=b\] Now you have a slope and y-intercept, so you have enough for the formula, so plug in your m and b values, giving you y=-6+57. Be sure to check your answer by plugging in the x value into this formula, and see if you get the same y value that is given.